Shift report 20000730 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Jul 30 2000 - 01:12:26 EDT

  • Next message: http account: "Shift report 20000730 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Jul 30 01:12:26 2000, Shift: 16:00-24:00
    Supervisor: KO, with: 
    Log book#3 pages 108 - 111
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine status: a few stores at injection energy,
       few attempts to accelerate, always beam lost during
       acceleration.
    
    2. Note from FV:
       - The new TPC pad monitor from Alv is now available
         in the operator account. Run "tpc3d", open a
         raw data from from "File" menu, select the TPC
         to look at.
       - recent data shows that T2 still does not work well.
    
    3. The latest T3, T4 and T5 high voltage settings were
       put into the HV spreadsheet and into the "BRAHMS
       Experiment Information" binder in the C.H. Please
       see the main log book for instructions. In the nutshell,
       the voltages on 2 sections of T3 should be turned on
       when we have beam.
    
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Jul 30 2000 - 01:16:05 EDT